OpenStudy (anonymous):

Type the equation of the given line in standard form. The line with a slope (-1/3) of and b of 6. Would I just substitute that into y=mx+b? giving me y=(-1/3)x+6. And then to get to standard form I would put the constants on one side giving me (-1/3)+y=6. Would that be right? Thanks ! :D

OpenStudy (hba):

did u try to solve it

OpenStudy (anonymous):

I don't have to solve it, just put it in standard form.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yep! standard form: \(ax^2+bx+c=0\) Can be found from simply re-arranging: y=mx+b y=(-1/3)x+6 Correct! Now when y = 0?

OpenStudy (anonymous):

You don't have any x^2 term, it's just something x and a constant :-)

OpenStudy (anonymous):

Oh okay, thanks :)

OpenStudy (anonymous):

np! ^_^